# A Gaussian smooth transition vector autoregressive model: An application to the macroeconomic effects of severe weather shocks

### v2

@@ Appendix B Monte Carlo algorithm for estimating the GIRF
    Repeat Steps 2-5 $R_{1}$ times and calculate the sample mean of $y_{t+h}^{(i)}(\delta_{j},\boldsymbol{y}_{t-1})-y_{t+n}^{(i)}(\boldsymbol{y}_{t-1})$ for $h=0,1,...,H$ to obtain an estimate of the GIRF$(h,\delta_{j},\boldsymbol{y}_{t-1})$.
  8. 7.
    Repeat Steps 1-6 $R_{2}$ times to obtain estimates of GIRF$(h,\delta_{j},\boldsymbol{y}_{t-1})$ with different starting values $\boldsymbol{y}_{t-1}$ generated from the distribution $G$. Then, take the sample mean and sample quantiles over the estimates to obtain point estimate and confidence intervals for the GIRF with random initial value.
Notice that if a fixed initial value $\boldsymbol{y}_{t-1}$ is used, Step 7 is redundant.
Appendix C Details on the empirical application
C.1 Model selection and adequacy of the selected model
@@
The maximum likelihood (ML) estimation of the models, residual diagnostics, and estimation of the generalized impulse response functions are carried out with the R package sstvars (Virolainen, 2024) in which the methods introduced in this paper have been implemented. The R package sstvars also contains the dataset studied in the empirical application to facilitate the reproduction of our results.
@@
We start model selection with a preliminary examination of the sample partial autocorrelation functions (PACF) of the series, which are presented in Figure C.1 for the first $24$ lags. Based on the breaks in the PACFs, an autoregressive order $p=4$ could be reasonable for parsimonious model, and an order smaller than $p=2$ would be clearly insufficient. Therefore, we estimate two-regime GSTVAR models with $p=1,...,5$ and find that the AIC is minimized by the order $p=4$, whereas BIC and HQIC are minimized by the order $p=2$. The values of the information criteria are presented in Table 1 (of the paper) together with the maximized log-likelihoods (divided by the number of observations).
@@
Figure C.1: Partial autocorrelation functions for the first $24$ lags estimated for the monthly U.S. series covering the period from 1961:1 to 2022:3.
To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the residuals for the first $24$ lags in Figure C.2. As the figure shows, there is not much autocorrelation left in the residuals, but somewhat large correlation coefficient (CC) sticks out at the lag seven in the ACF of the interest rate variable. Moreover, there are somewhat large CCs at the lag zero also in the CCFs between CPI and GDP. Nevertheless, the fitted models seems to capture the autocorrelation structure of the data reasonably well.
In order to study the model’s adequacy to capture the conditional heteroskedasticity of the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure C.3. The figure shows that there are several large CCs in the ACFs of the squared standardized residual of GDP and CPI. Otherwise, the CCs are not particularly large. That is, some of the conditional heteroskedasticity of GDP and CPI is not adequately captured, but in our view, the inadequacies are not very severe.
The series of standardized residuals, presented in the top panels of Figure C.4, also show some remaining heteroskedasticity and several outliers. There is a particularly large negative residual of the GDP growth rate in the beginning of the COVID-19 crisis, when the lockdown of the economy caused a substantial drop in the GDP. However, since the COVID-19 drop was caused by an exceptionally large exogenous shock, a large residual is expected for a correctly specified model. Finally, the normal quantile-quantile-plots, presented in the bottom panel of Figure 1, show that the marginal distribution of ACI is captured well. The marginal distributions of the residuals of the other variables, however, display excess kurtosis, but they are quite symmetric. Nonetheless, the overall adequacy of the model seems reasonable enough for impulse response analysis.
Finally, we assess whether the regime shifts can be attributed solely to changes in the level or volatility of the series, or whether the data exhibit time-variation also in the autoregressive dynamics. Since our GSTVAR model is ergodic stationary, the results of consistency and conventional limiting distribution apply to the ML estimator under conventional high level conditions, such that standard likelihood based tests are available. Hence, to check if the regime shifts are driven only by changes in the volatility regime, we test whether the intercepts and AR matrices are identical across the regimes (i.e., time-invariant). The Wald test for the null of constancy of intercepts and AR matrices produces a $p$-value smaller than $10^{-7}$, indicating that the regime-shifts are not solely driven by changes in the volatility regime. The Wald test for the constancy of the AR matrices alone produces a $p$-value smaller than $10^{-5}$. Therefore, we conclude in favor of regime shifts in the level, volatility, as well as the autoregressive dynamics of the series.
@@
Figure C.2: Auto- and crosscorrelation functions of the residuals of the fitted two-regime GSTVAR $p=4$ model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure C.3: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime fourth-order GSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure C.4: Standardized residual time series and normal quantile-quantile-plots of the fitted two-regime fourth-order GSTVAR model.
C.2 Robustness checks
As a robustness check, we consider the subsample period from 1963:4 to 2019:5 studied by Kim et al. (2022). The ACI and interest rate variable series are similar to Kim et al. (2022), but our data differs from them in two main aspects. First, instead of using unemployment rate and year-on-year industrial production index as a measure of real economic activity, we use the monthly GDP growth rate described in Section 4. Second, instead of using year-on-year CPI inflation rate, we use the monthly CPI inflation growth rate.
Figure C.5: Estimated transition weights of Regime 1 for the two-regime fourth-order GSTVAR model fitted to the monthly U.S. data covering the period from 1963:4 to 2019:5. The shaded areas indicate the period of NBER based U.S. recessions.
@@ C.2 Robustness checks
We fit two-regime GSTVAR models the subsample and select the autoregressive order $p=4$ based on information criteria and preliminary examinations of partial autocorrelation functions of the series (not shown). The estimated transition weights of our two-regime fourth-order GSTVAR model are presented Figure C.5 with the shaded areas indicating the periods of NBER based on U.S. recessions. As the transition weights show, the regimes are quite similar to the benchmark specification: Regime 1 mainly dominates in the volatile period of 1960’s, 1970’s, and 1980’s, but obtains obtains large weights also in the 1990’s recession and the Financial crisis. Conversely, Regime 2 dominates when the first one does not.
@@
The GIRFs estimated for the model are presented in Figure C.6, which is similar to Figure 2 of the main paper. Responses to shocks of different sign and size seem similar, so only the responses to a positive one-standard-error shock are reported. The GIRFs are mostly similar to our main specification utilizing the full sample period from 1961:1 to 2022:3. In both of the regimes, a positive ACI shock decreases GDP and consumer prices, but the effects are stronger in Regime 1. The GIRFs are close to linear in Regime 1, while there is variation in the GIRFs depending on the initial values in Regime 2. A notable difference to our main specification is that a positive ACI shock increases the interest rate in Regime 1 but decreases it in Regime 2.
@@
Figure C.6: Generalized impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead, based on the data from from 1963:4 to 2019:5. From the top to bottom, the responses of the ACI index, GDP growth rate, inflation growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in the rows. The left column shows the responses to a positive one-standard-error shocks with initial values generated from the stationary distribution of Regime 1. The right column shows the corresponding responses in Regime 2. All GIRFs have been scaled so that instantaneous effect of the ACI index is $0.3$. The $95\%$ confidence intervals that reflect uncertainty about the initial state within the given regime are depicted with shaded grey areas. The results are based on 2500 draws of initial values (see Appendix B for details).

### v3

@@ Appendix B Monte Carlo algorithm for estimating the GIRF
    Repeat Steps 14 $R_{1}$ times and calculate the sample mean of $y_{t+h}^{(i)}(\delta_{j},\boldsymbol{y}_{t-1})-y_{t+n}^{(i)}(\boldsymbol{y}_{t-1})$ for $h=0,1,...,H$ to obtain an estimate of the GIRF$(h,\delta_{j},\boldsymbol{y}_{t-1})$.
To estimate the GIRF for a number of histories (and possibly signs and sizes of the shock), repeat the above steps for each of the initial values (and signs and sizes of the shocks), either taking them from the set $S$ or generating them from the distribution $G$. The sign and size $\delta_{j}$ of the shock can vary depending on the initial value or it can be fixed throughout. To obtain summary statistics on the distribution of the GIRFs over the initial values (or signs or sizes of the shock), the sample mean or some sample quantiles over the GIRFs can be calculated.
Appendix C Comparison to related models
In this appendix, we compare the GSTVAR model to a number of alternative closely related models popular in the literature, namely the GMVAR, LSTVAR, TVAR, MS-VAR, and TVP-VAR models.
Although the transition weight function of the GSTVAR model is similar to the mixing weights of the GMVAR model, it differs from that model in other ways. Specifically, the GMVAR model generates each observation from a randomly selected distinct regime, while the GSTVAR model allows for smooth transitions between the regimes and thereby facilitates gradual shifts in the dynamics. This is the case because in the GMVAR model, the transition weights $\alpha_{m,t}$ are replaced by unobservable regime variables $s_{m,t}$ in Equations (2.1)–(2.3). These regime variables $s_{m,t}$, $m=1,...,M$, are such that at each $t$, exactly one of them takes the value one, whereas the rest take the value zero according to the probabilities $P(s_{m,t}=1|\mathcal{F}_{t-1})=\alpha_{m,t}$. Because the regime that generates each observation is thus unknown in the GMVAR model, the (structural) errors cannot, in general, be recovered from the data, while they are readily available in the GSTVAR model. This is particularly useful in structural analysis because it enables to compute the historical decomposition and conducting counterfactual analyses, and the recovered structural shocks also facilitate computing GIRFs that reflect the properties of the data more precisely (see Sections 2.2 and 4).
$\alpha_{1,t}=1-\alpha_{2,t},\ \ \alpha_{2,t}=[1+\exp\{-\gamma(y_{i,t-j}-c)\}]^{-1},$ (C.1)
where $y_{i,t-j}$ is the $j$th lagged observation $(j\in\{1,....,p\})$ of the $i$th variable $(i\in\{1,....,d\})$, $c\in\mathbb{R}$ is a location parameter, and $\gamma>0$ is a scale parameter. The location parameter $c$ determines the mid-point of the transition function, i.e., the value of the switching variable when the weights are equal. The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $y_{it-j}$. This simple setup can be generalized in various directions (see, e.g., Teräsvirta and Yang, 2014). It is also possible to use exogenous or deterministic switching variables instead of a lagged endogenous variable in the logistic transition weights (C.1). However, in these cases, it may be difficult to determine whether the model is stationary, and potentially important dynamics are ignored, as the shocks are not allowed affect the transition weights, contradicting the premise that economic agents may alter their behavior when facing changing economic conditions.
The logistic transition weights have the advantage that the regimes typically have clear interpretations and the switching variable can be defined to address some specific aspect of the empirical problem. For example, when using the lagged GDP growth as the switching variable, the regimes can be interpreted as recession and expansion. In comparison, the transition weights of the GSTVAR model depend on the full distribution of the preceding $p$ observations and can thereby capture more complicated switching-dynamics, which facilitates associating the regimes with various aspects of the phenomenon being modelled. As the transition weights of the GSTVAR model are not affected by the choice of switching variables but determined based on the weighted relative likelihoods of the regimes, they are more clearly data-driven and their own evolution is interesting as such. Nonetheless, since the resulting regimes are statistically determined, it is not always clear how to interpret them in terms of a given application.
$\alpha_{m,t}=I(c_{m-1}<s_{t}<c_{m}),\ \ m=1,...,M,$ (C.2)
$P(s_{m,t}=1|s_{n,t-1}=1)=p_{mn}\ \text{for all}\ t\ \text{and}\ m,n=1,...,M,$ (C.3)
that satisfy $\sum_{n=1}^{M}p_{mn}=1$ for all $m$. In other words, the regime that generates each observation is selected randomly at each $t$ according to the probabilities $p_{mn}$, $m,n=1,...,M$. While the MS-VAR model can flexibly accommodate unobservable discrete regime-switches, unlike our GSTVAR model, it is unable to capture gradual shifts in the dynamics of the data. Moreover, in the same way as in the GMVAR model, it is not possible to recover structural shocks from the fitted model because the regime that generates each observation is not observed.
Appendix D A Monte Carlo study
To assess the properties of the ML estimator, we perform a small scale Monte Carlo study. Because the estimation of our GSTVAR model is computationally demanding, we assume a simple setup with two variables ($d=2$), two regimes ($M=2$), and autoregressive order one ($p=1$) in model (2.1)–(2.3). Two different sets of values of the parameter $\theta=(\phi_{1,0},\phi_{2,0},\text{vec}(A_{1,1}),\text{vec}(A_{2,1}),\text{vech}(\Omega_{1}),\text{vech}(\Omega_{2}),\alpha_{1})$ $(19\times 1)$ are considered. We refer to the resulting models as as Model 1 and Model 2. In each specification, we use the same values of the covariance matrix and the transition weight parameter, $\text{vech}(\Omega_{1})=(0.50,0.20,0.30)$, $\text{vech}(\Omega_{2})=(0.80,-0.20,0.50)$, and $\alpha_{1}=0.70$, but the intercepts and AR matrices differ. Hence, we we are able to learn about the importance of the dynamics of the model for the finite sample properties of the estimator.
The employed values of $\phi_{1,0},\phi_{2,0},\text{vec}(A_{1,1}),\text{vec}(A_{2,1})$ are presented in Table D.1 for Model 1 and Model 2, along with the unconditional means as well as the modulus of the eigenvalues of the companion form AR matrices of the regimes. The latter indicate that Model 1 does not exhibit very high degree of persistence, whereas the AR matrices of Model 2 are somewhat close the boundary of the stationarity region. There is a reasonable amount of variation in the unconditional means across the regimes in both specifications, with less (more) variation for the first (second) component series in Model 1, and vice versa in Model 2.
Regime 1 Regime 2
$\phi_{1,0}$ $A_{1,1}$ $\mu_{1}$ $||\boldsymbol{A}_{1}||$ $\phi_{2,0}$ $A_{2,1}$ $\mu_{2}$ $||\boldsymbol{A}_{2}||$
[…]
Model 2 $0.00$ $0.80$ $-0.55$ $-2.29$ $0.97$ $0.50$ $-0.99$ $-0.20$ $-0.19$ $0.96$
[…]
Table D.1: The parameters values of $\phi_{1,0},\phi_{2,0},A_{1,1}$, and $A_{2,1}$ used for Model 1 and Model 2 in the Monte Carlo study. Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m,0}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
$T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
[…]
$\phi_{2,0}$ $-0.35$ $(1.10)$ $-0.18$ $(0.71)$ $-0.08$ $(0.48)$ $-0.04$ $(0.32)$ $-0.02$ $(0.15)$
[…]
Table D.2: Results from the Monte Carlo study on the performance of the ML estimator based on $500$ repetitions. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses.
For both model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$. Then, we fit the first-order two-regime GSTVAR model to each of the generated samples based on eight estimation rounds of the two-phase procedure proposed in Section 3. To estimate the bias of the ML estimator, we calculate the average estimation errors $\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$ over the $500$ Monte Carlo repetitions, and to estimate the estimation accuracy, we calculate the standard deviations of the estimates.
The results in Table D.2 show that for all the parameters, the bias decreases and estimation accuracy increases substantially as the sample size increases, in line with the consistency of the ML estimator. There is some but mostly small bias when the sample size is small, and particularly the intercept parameters are inaccurately estimated. For most of the parameters, the estimation accuracy seems quite reasonable already in samples of length $500$, however. In the samples of length $2000$ and $10000$, the bias has practically disappears and estimation accuracy is very high for all but some of the intercept parameters, and also the intercept parameters are reasonably well estimated. Finally, there does not appear to be notable differences in estimation accuracy between the two model specifications.
Appendix E Details on the empirical application
E.1 Model selection and adequacy of the selected model
@@
Maximum likelihood (ML) estimation of the models, residual diagnostics, and estimation of the generalized impulse response functions are carried out with the R package sstvars (Virolainen, 2025a) in which the methods introduced in this paper have been implemented. It also contains the dataset studied in the empirical application to facilitate the reproduction of our results.
@@
We start model selection with a preliminary examination of the sample partial autocorrelation functions (PACF) of the series, which are presented in Figure E.1 for the first $24$ lags. Based on the breaks in the PACFs, an autoregressive order $p=4$ could be reasonable for parsimonious model, and an order smaller than $p=2$ would be clearly insufficient. Therefore, we estimate two-regime GSTVAR models with $p=1,...,5$ and find that the AIC is minimized by the order $p=4$, whereas BIC and HQIC are minimized by the order $p=2$. The values of the information criteria are presented in Table 1 (of the paper) together with the maximized log-likelihoods (divided by the number of observations).
@@
Figure E.1: Partial autocorrelation functions for the first $24$ lags estimated for the monthly U.S. series covering the period from 1961:1 to 2022:3.
@@
Figure E.2: Auto- and crosscorrelation functions of the residuals of the fitted two-regime GSTVAR $p=4$ model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure E.3: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime fourth-order GSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure E.4: Standardized residual time series and normal quantile-quantile-plots of the fitted two-regime fourth-order GSTVAR model.
To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the residuals for the first $24$ lags in Figure E.2. As the figure shows, there is not much autocorrelation left in the residuals, but somewhat large correlation coefficient (CC) sticks out at the lag seven in the ACF of the interest rate variable. Moreover, there are somewhat large CCs at the lag zero also in the CCFs between CPI and GDP. Nevertheless, the fitted models seems to capture the autocorrelation structure of the data reasonably well.
In order to study the model’s adequacy to capture the conditional heteroskedasticity of the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure E.3. The figure shows that there are several large CCs in the ACFs of the squared standardized residual of GDP and CPI. Otherwise, the CCs are not particularly large. That is, some of the conditional heteroskedasticity of GDP and CPI is not adequately captured, but in our view, the inadequacies are not very severe.
The series of standardized residuals, presented in the top panels of Figure E.4, also show some remaining heteroskedasticity and several outliers. There is a particularly large negative residual of the GDP growth rate in the beginning of the COVID-19 crisis, when the lockdown of the economy caused a substantial drop in the GDP. However, since the COVID-19 drop was caused by an exceptionally large exogenous shock, a large residual is expected for a correctly specified model. Finally, the normal quantile-quantile-plots, presented in the bottom panel of Figure E.4, show that the marginal distribution of ACI is captured well. The marginal distributions of the residuals of the other variables, however, display excess kurtosis, but they are quite symmetric. Nonetheless, the overall adequacy of the model seems reasonable enough for impulse response analysis.
In our model, the covariance matrix of the errors as well as the autoregressive matrices and the intercept term can vary between the regimes. However, it is possible that the regime shifts are attributed solely to changes in the level or volatility of the series and there is no time-variation in the autoregressive dynamics. Since our GSTVAR model is ergodic stationary, the results of consistency and conventional limiting distribution apply to the ML estimator under conventional high level conditions and standard likelihood based tests can be used to check if the regime shifts are driven only by changes in the volatility regime. The Wald test for the null of constancy of intercepts and AR matrices produces a $p$-value smaller than $10^{-7}$, indicating that the regime-shifts are not solely driven by changes in the volatility regime. The Wald test for the constancy of the AR matrices alone produces a $p$-value smaller than $10^{-5}$. Therefore, we conclude in favor of regime shifts in the level, volatility, as well as the autoregressive dynamics of the series.
E.2 Robustness checks
E.2.1 The GSTVAR model fitted to a sub-sample
@@
As the first robustness check, we consider the subsample period from 1963:4 to 2019:5 studied by Kim et al. (in press). [^3] The ACI and interest rate variable series are similar to theirs, but otherwise our data differ in two main aspects. First, instead of using unemployment rate and year-on-year industrial production index as a measure of real economic activity, we use the monthly GDP growth rate described in Section 4. Second, instead of using year-on-year CPI inflation rate, we use the monthly CPI inflation growth rate. We fit two-regime GSTVAR models the subsample and select the autoregressive order $p=4$ based on information criteria and preliminary examinations of partial autocorrelation functions of the series (not shown). The estimated transition weights of our two-regime fourth-order GSTVAR model are presented Figure E.5 with the shaded areas indicating the periods of NBER based on U.S. recessions. As the transition weights show, the regimes are quite similar to the benchmark specification: Regime 1 mainly dominates in the volatile period of 1960s, 1970s, and 1980s, but obtains obtains large weights also in the 1990s recession and the Financial crisis. Conversely, Regime 2 dominates in the less volatile periods.
Figure E.5: Estimated transition weights of Regime 1 for the two-regime fourth-order GSTVAR model fitted to the monthly U.S. data covering the period from 1963:4 to 2019:5. The shaded areas indicate the period of NBER based U.S. recessions.
@@
The GIRFs estimated for the model are presented in Figure E.6, which is similar to Figure 2 of the paper. The GIRFs are also mostly similar to our main specification utilizing the full sample period from 1961:1 to 2022:3. In both of the regimes, a positive ACI shock decreases GDP and consumer prices, but the effects are stronger in Regime 1. A notable difference to our main specification is that a positive ACI shock increases the interest rate in Regime 1, but decreases it in Regime 2.
@@
Figure E.6: Generalized impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead, based on the data from from 1963:4 to 2019:5. From the top to bottom, the responses of ACI, GDP growth rate, CPI growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in the rows. The left and right columns show the responses conditional on Regime 1 and Regime 2 dominating when the shock arrives, respectively. All GIRFs have been scaled so that instantaneous effect of the ACI index is $0.3$. The results are based on $2500$ Monte Carlo repetitions for each history (see Appendix B for details).
E.2.2 The GMVAR model and a linear SVAR model
For comparison, we also fit the two-regime fourth-order GMVAR model (Kalliovirta et al., 2016) as well as the recursively identified linear Gaussian SVAR model (with autoregressive order $p=9$ based on the AIC) to the full sample period. The estimated mixing weight of the GMVAR model, presented in Figure E.7 (black dashed line), is very similar to the estimated transition weights of our GSTVAR model. Hence, the regimes of the GMVAR model and our GSTVAR model have similar interpretations.
Figure E.8 presents the GIRFs to a positive one-standard-error ACI shock in the GMVAR model (black dashed line) with the histories generated from the stationary distribution of each regime. Only the average GIRFs over the histories are depicted for clarity. For easy comparison, also the average GIRFs produced in a similar manner for our GSTVAR model are depicted (grey solid line). The GIRFs based on the GMVAR model are quite similar to those of our GSTVAR model in the sense that in both regimes, GDP and the interest rate decrease and there is a short-term decrease in inflation rate. Like in our GSTVAR model, the responses of GDP, inflation rate, and interest rate are clearly stronger in the more turbulent Regime 1, and the responses of the mixing weights are very weak in both regimes.
To study possible asymmetries in the effects of the ACI shock with respect to its sign and size, we calculate the GIRFs also for a negative one-standard-error ACI shock as well as for positive and negative two-standard-error ACI shocks. However, marked asymmetries in neither the GMVAR nor the GSTVAR model can be found. Therefore, to save space, the GIRFs for negative or two-standard-error ACI shock are not presented.
Figure E.8 also presents the impulse response functions (IRF) to the ACI shock in the linear SVAR model (black dotted line, depicted in both of the columns of Figure E.8). Like in the GSTVAR model, the IRFs show a decrease in GDP, CPI and interest rate in response to a positive ACI shock, but the linear model is, of course, incapable of capturing the differences across the regimes.
Figure E.7: Estimated mixing weights of Regime 1 of the two-regime fourth-order GMVAR model fitted to the full sample period (black dashed line) together with the estimated transition weights of Regime 1 of our GSTVAR model (grey solid line).
Figure E.8: The (generalized) impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead in the GSTVAR model (grey solid line), the GMVAR model (black dashed line), and the linear Gaussian SVAR model (black dotted line). From top to bottom, the responses of ACI, GDP growth rate, CPI growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in the rows. The left and right columns show the responses to a positive one-standard-error shock with histories generated from the stationary distribution of Regime 1 and Regime 2, respectively. All impulse responses have been scaled so that the instantaneous change in the ACI index is $0.3$. The GIRFs are based on $2500$ draws of initial values (for details of the algorithm, see Appendix B). For the linear SVAR, the conventional IRFs are depicted in both columns.